OpenStudy (saifoo.khan):

Simple question:

OpenStudy (saifoo.khan):

im getting 225%

terenzreignz (terenzreignz):

Likewise.

OpenStudy (saifoo.khan):

Hey, ain't it gonna be like 225% of the whole. which means 225% - 100% = 125% ?

OpenStudy (shubhamsrg):

why not 125% ?

OpenStudy (shubhamsrg):

yes,,i got 125%

OpenStudy (saifoo.khan):

And the logic is same what i just gave right? @shubhamsrg

terenzreignz (terenzreignz):

Oh right, lol, I keep overlooking these details... you got me there ;)

OpenStudy (callisto):

Original diameter = d radius = 0.5d New diameter = 1.5d new radius = 0.75d Area change = (0.75^2 - 0.5^2)/0.5^2 x 100% = 125%

OpenStudy (shubhamsrg):

not exactly.. i used orthodox method. let diameter be x so area = pi(x^2)/4 new diameter = 3x/2 so new area = pi(9x^4)/4 so % inc in area = (pi(9x^4)/4 - pi(x^2)/4 )/pi(x^2)/4 * 100 which directly gives 125%

OpenStudy (callisto):

Area change = (pi 0.75^2 d - pi 0.5^2 d)/(pi 0.5^2 d) x 100% = 125% Sorry I was lazy...

OpenStudy (shubhamsrg):

exactly what @Callisto said! :)

OpenStudy (saifoo.khan):

Thanks! =D

OpenStudy (saifoo.khan):

Unfortunately i can give only 1 medal. :/

OpenStudy (callisto):

To @shubhamsrg !!

OpenStudy (shubhamsrg):

fortunately,,you gave it to the right person! :D ;)

OpenStudy (shubhamsrg):

medal has already been given miss!! nevermind! ;)

OpenStudy (saifoo.khan):

Haha, I owe you one! Will give you somwhere else.. Don't worry man! :D @shubhamsrg

OpenStudy (shubhamsrg):

heheheh..its a deal then! :D

OpenStudy (saifoo.khan):

;)